OpenStudy (anonymous):

find the dimensions of a rectangular garden with perimeter 62 yd and area 220 yd^2

OpenStudy (anonymous):

Its a system of equation. Perimeter = 2(length) + 2(width) Area = (length) * (width)

OpenStudy (anonymous):

set the Perimeter = to 62 and the area = to 220

OpenStudy (anonymous):

i really just need a step by step to get the answer i dont get it whatsoever

OpenStudy (anonymous):

b*h = 220 2b+2h = 62 b = 31-h substitution (31-h)*h = 220 h^2-31h+220 = 0 quadratic h = 11, 20 b = 20, 11

OpenStudy (anonymous):

so dimensions are 11x20